# Do Multimodal LLMs See Sentiment?

### v1

@@ Abstract
Index Terms:
Visual Sentiment Analysis, Multimodal Large Language Models, Computational Social Systems, Social Media Analytics.
@@ I Introduction
  • We release the generated image descriptions for two datasets, PerceptSent and DeepSent, along with our fine-tuned models to ease future research and reproducibility [^1] Will be provided after journal publication..
@@ II Related work
Textual Sentiment Analysis. The field of textual sentiment analysis has evolved significantly from early lexicon-based approaches [16, 17] to sophisticated deep learning models. Traditional methods rely on sentiment dictionaries and rule-based systems [18, 19], whereas modern approaches leverage neural architectures, such as BERT-based models [20] and graph neural networks [21, 22]. A recent work [23] demonstrates that transformers can capture aspect-level sentiment through double-view graph representations. The success of textual analysis has led to its application in multimodal contexts, such as video transcriptions [24]. However, challenges remain when analyzing platforms like Instagram where visual content dominates and textual context is sparse [25].
@@
Multimodal Approaches. Combining visual and textual modalities has emerged as a powerful paradigm for sentiment analysis [30]. For low-resource languages, [31] demonstrates the effectiveness of late fusion strategies, whereas [32] proposes interaction networks to capture cross-modal dependencies. Recent work has introduced sophisticated architectures such as the multi-stage perception model in [33] and the dual-perspective fusion network in [34], which better align visual and textual representations. These methods show that joint modeling of modalities can overcome limitations of unimodal approaches, particularly for ambiguous or context-dependent sentiment expressions. These advances in multimodal fusion models set the stage for a new paradigm, where vision-language understanding is embedded directly into large-scale language models.
@@
This paper builds upon these advances by systematically evaluating how MLLMs can address key challenges in visual sentiment analysis, particularly in handling subjective interpretations and bridging the semantic gap between visual content and affective responses. Our work extends current approaches by comparing different architectures and strategies for leveraging image-to-text transformations in sentiment prediction.
@@ III Datasets and Problem Formulation
In the present study, we use two datasets: PerceptSent [5] and DeepSent [26]. The former is a publicly available [^2] https://github.com/ceslop84/perceptsent. collection of 5,000 images, most (92.7%) of which depict outdoor scenes. The images have been sourced from Instagram, Flickr, and NYC311333https://portal.311.nyc.gov.. Designed specifically for sentiment analysis, the dataset has been fully annotated by human evaluators using mainly Amazon Mechanical Turk [^4] https://www.mturk.com. (AMT). The latter consists of 1,269 images collected from X (old Twitter), from indoor and outdoor scenes, which have also been labeled through AMT crowd-sourcing platform.
To align with prior work and facilitate reproducibility, we adopt the problem formulation for sentiment analysis proposed by Lopes [5]. Denoted here as $\langle{\sigma_{l}},{P}_{{C}}\rangle$, we have $C$ as the number of sentiment categories considered in a problem instance $P_{C}$, and $l$ as a threshold for filtering out divergent evaluators opinions, allowing us to define different types of sentiment dominance for filter $\sigma$.
@@ IV-A Multimodal Large Language Models (MLLMs)
To ensure effective task alignment, we have experimented with the aforementioned MLLMs with various prompt formulations, including those generated by the models themselves through prompt engineering and self-instruction. After qualitative and quantitative comparisons on a representative subset of samples, we selected the most effective prompt for each task, balancing clarity, consistency, and model response accuracy:
@@ IV-B Text-only Large Language Models (LLMs)
When adjusting the weights in Task 2a and Task 2b, the training loop iteratively updates model parameters over several epochs and incorporates early stopping based on validation F1-score improvements to prevent overfitting. The instances used as LLM inputs are textual descriptions provided by MLLMs. Such descriptions are framed as text generation tasks using sentiment-aware prompts adapted to the number of classes in each configuration $P_{C}$, where $C\in\{2,3,5\}$. Each prompt instructs the model to determine the sentiment of a given description by selecting from a predefined set of labels. For example, in the three-class setting $P_{3}$, the prompt is “What is the sentiment of this description? Please choose an answer from {“Positive”: 2, “Negative”: 0, “Neutral”: 1}”. Then they are tokenized and go through data-loaders with dynamic batching and shuffling for robust training.
@@ V-A Direct Sentiment Classification using MLLMs
As shown in the third column of Table I (Task 1), the lowest $F$-scores for GPT (OAI) and DeepSeek occur under more challenging conditions, characterized by low annotator agreement, while higher scores are observed as annotator consensus increases. The same trend is observed with respect to the number of target classes; lower scores are obtained under harsher conditions, i.e., as the number of classes increases. GPT (OAI) achieves higher F-scores than DeepSeek, with relative gains [^5] For example, the relative $F$-score gain of GPT (OAI) regarding DeepSeek (DS) is calculated as (OAI - DS)/DS of 2% to 67%. The improvement is statistically significant (paired t-test, $p<0.05$) in three cases; however, in the $\langle\sigma_{5},P_{3}\rangle$ configuration, the one with the lowest difference, this value is not significant ($p=0.194$).
In contrast, GPT (OS) fails to produce valid classification outputs for Task 1. As illustrated in Fig. 2, even after extensive prompt engineering — including explicit instructions such as “Select a single sentiment class for this image from the $\{\mathcal{L}\}$ list”, where $\{\mathcal{L}\}$ refers to the list of sentiment polarities for a specific problem setup, or “Do not describe this image; just choose one sentiment polarity from the $\{\mathcal{L}\}$ list” — the model consistently generates explanatory or ambiguous responses that attempt to justify the set $\{\mathcal{L}\}$ of categories, rather than selecting a single label, or combining multiple sentiment cues (e.g., “Positive, but with some neutral elements”), making it difficult to unambiguously assign a single label and complicating automatic parsing. This behavior contrasts with the other MLLMs, which produce discrete and task-aligned outputs, as also shown in Fig. 2.
@@
Refer to caption Problem setup: P3 Dominance threshold: $\sigma=3$ Labeled as: Negative Task 1 GPT (OS) Positive: The image shows a group of people walking on the sidewalk, which suggests that they are out and about, enjoying their day or going about their daily business. The bright sunlight and clear blue sky add to the positive sentiment of the image. Neutral: The image is a reflection in a round mirror, making it difficult to determine the sentiment of the scene beyond the mirror … Negative: It is not possible to determine the sentiment of the image due to the reflection in the round mirror. The reflection obscures the view … GPT (OAI): Negative DeepSeek: Neutral 10121Evaluator’s votes
@@ V-B Pre-trained LLMs: classifying MLLMs image descriptions
V-C Fine-tuning LLMs: classifying MLLMs descriptions
As shown in Table I, fully fine-tuning the LLMs (Task 2b) yields the highest overall $F$-scores, a consistent trend across all configurations. Relative gains over Task 2a range from [11.59%, 50.87%] for BART, [5.74%, 32.20%] for MBERT, and [4.09%, 102.87%] for LLAMA, across all MLLMs, evaluators’ agreement thresholds, and number of sentiment classes.
@@ V-C Fine-tuning LLMs: classifying MLLMs descriptions
Fig. 4 showcases four examples from the PerceptSent dataset alongside sentiment predictions from all considered models in both Task 2a and Task 2b. Figs. 4(a,b) present image samples with strong consensus ($\sigma_{5}$) among human evaluators. In Fig. 4(a), the image has received unanimous positive sentiment annotations from all five annotators. After fine-tuning, all MLLM+LLM combinations correctly predict the positive sentiment based on generated descriptions referencing visual elements such as a body of water, a beach, and sparkling highlights. Fig. 4(b) similarly shows high agreement, with all MLLM+LLM combinations (pre-trained and fine-tuned) accurately identifying the target negative sentiment. The image description results often highlight features such as an underground tunnel, a graffiti-covered entrance, and a gritty atmosphere, with GPT (OAI) explicitly mentioning a sense of decay. Interestingly, both GPT (OAI) and DeepSeek can even recognize and transcribe the “OHNO” graffiti above the entrance. As shown in Table I, these two cases exemplify the best observed performance, where high evaluator agreement ($\sigma_{5}$) correlates with elevated $F$-scores. In contrast, under lower agreement ($\sigma_{3}$), some issues occur. Although in Fig. 4(c) most of the pre-trained models correctly classify slightly negative sentiment, some BART predictions shift toward wrong sentiment after fine-tuning. This change may result from the influence of image descriptions generated by GPT (OAI), such as “The atmosphere looks calm and residential, typical of a suburban area,” which potentially moderates the models’ sentiment evaluations. Nevertheless, fine-tuning improvements occur, particularly for GPT (OS)+LLAMA and DeepSeek+MBERT. In the case illustrated by Fig.4(d), none of the MLLM+LLM combinations succeed in predicting the correct negative sentiment label after fine-tuning (Task 2b). This fact appears to be driven by limitations in the scene descriptions generated by the MLLMs. Specifically, none of the models could identify salient visual features indicative of a negative context — such as the pile of trash on the left or the overall untidy condition of the street. Instead, the generated image descriptions emphasize neutral or positive aspects, including references to a city street with parked cars, a calm atmosphere, and the presence of a few pedestrians. Such framing has likely influenced the LLMs to interpret the scene as neutral. Notably, LLAMA in Task 2a predicts the correct negative sentiment; however, with low $F$-scores ($30.3\%,46.6\%]$) for $\langle\sigma_{3},P_{3}\rangle$, this does not indicate consistent performance. After fine-tuning, the prediction shifted to neutral, confirming that the generated descriptions failed to capture the negative elements identified by human evaluators.
@@
Refer to caption Problem setup: P5 Dominance threshold: $\sigma_{5}$ Labeled as: Positive GPT (OS) GPT (OAI) DeepSeek BART Task 2a: Positive Positive Slightly Pos. Task 2b: Positive Positive Positive MBERT Task 2a: Slightly Pos. Positive Positive Task 2b: Positive Positive Positive LLAMA Task 2a: Positive Positive Positive Task 2b: Positive Positive Positive 50000Evaluator’s votes(a)
@@
Refer to caption Problem setup: P3 Dominance threshold: $\sigma_{5}$ Labeled as: Negative GPT (OS) GPT (OAI) DeepSeek BART Task 2a: Negative Negative Negative Task 2b: Negative Negative Negative MBERT Task 2a: Negative Negative Negative Task 2b: Negative Negative Negative LLAMA Task 2a: Negative Negative Negative Task 2b: Negative Negative Negative 00014Evaluator’s votes(b)
@@
Refer to caption Problem setup: P5 Dominance threshold: $\sigma_{3}$ Labeled as: Slightly Negative GPT (OS) GPT (OAI) DeepSeek BART Task 2a: Slightly Neg. Slightly Neg. Slightly Neg. Task 2b: Negative Neutral Slightly Neg. MBERT Task 2a: Slightly Neg. Slightly Neg. Negative Task 2b: Slightly Neg. Slightly Neg. Slightly Neg. LLAMA Task 2a: Positive Neutral Slightly Neg. Task 2b: Slightly Neg. Neutral Slightly Neg. 00131Evaluator’s votes(c)
@@
Refer to caption Problem setup: P3 Dominance threshold: $\sigma_{3}$ Labeled as: Negative GPT (OS) GPT (OAI) DeepSeek BART Task 2a: Positive Neutral Positive Task 2b: Neutral Neutral Positive MBERT Task 2a: Neutral Neutral Neutral Task 2b: Neutral Neutral Neutral LLAMA Task 2a: Negative Negative Negative Task 2b: Neutral Neutral Neutral 00212Evaluator’s votes(d)
@@ VI Conclusions and future work
Although we evaluated GPT-4o mini, MiniGPT-4 and DeepSeek-VL2-Tiny, our methodology is model-agnostic. The consistent performance patterns suggest newer MLLMs would yield further improvements while maintaining interpretability benefits. Notably, ModernBERT’s superior performance over larger models demonstrates that careful architecture selection and tuning can outweigh pure scale advantages. Moreover, despite GPT-4o mini achieving the best overall results, we note that open-weights (DeepSeek) or open source (GPT (OS)) alternatives can still be attractive compared to paid commercial models such as those from OpenAI. In that regard, we highlight that, due to GPU and time constraints, we employed the Tiny DeepSeek variant, and that larger models could achieve performance levels closer to that of GPT-based models. Thus, we believe a deeper exploration of open-weights and open source models remains an interesting road for future research.
@@
We identify, therefore, three key future directions: (1) exploring our pipeline to sensitive domains such as mental health monitoring, where sentiment analysis can provide nuanced signals to support user well-being, and in the detection of potentially harmful or distressing content in social media; (2) integrating xAI techniques to make the model’s reasoning process more interpretable, e.g., highlighting which regions in the image and which words in the generated description most influenced the sentiment classification; and (3) developing interfaces that let users correct ambiguous image descriptions to improve model performance.

### v2

# Multimodal LLMs See Sentiment

@@ I Introduction
  • We release the generated image descriptions for two datasets, PerceptSent and DeepSent, along with our fine-tuned models to ease future research and reproducibility [^1] The link will be public soon – contact the authors..
@@ II Related work
Textual Sentiment Analysis. The field of textual sentiment analysis has evolved significantly from early lexicon-based approaches [16, 17] to sophisticated deep learning models. Traditional methods rely on sentiment dictionaries and rule-based systems [18, 19], whereas modern approaches leverage neural architectures, such as BERT-based models [20] and graph neural networks [21, 22]. A recent work [23] demonstrates that transformers can capture aspect-level sentiment through double-view graph representations. The success of textual analysis has led to its application in multimodal contexts, such as video transcriptions [24]. However, challenges remain when analyzing platforms like Instagram where visual content dominates, and textual context is sparse [25].
@@
Multimodal Approaches. Combining visual and textual modalities has emerged as a powerful paradigm for sentiment analysis [30]. For low-resource languages, [31] demonstrates the effectiveness of late fusion strategies, whereas [32] proposes interaction networks to capture cross-modal dependencies. Recent work has introduced sophisticated architectures such as the multi-stage perception model [33] and the dual-perspective fusion network [34], which better align visual and textual representations. These methods show that joint modeling of modalities can address the limits of unimodal approaches, especially for ambiguous or context-dependent sentiment expressions. These advances in multimodal fusion models set the stage for a new paradigm, where vision-language understanding is embedded into large language models.
@@
This paper builds on these advances by evaluating how MLLMs address key challenges in visual sentiment analysis, especially in handling subjective interpretations and bridging the semantic gap between visual content and affective responses. Our work extends current approaches by comparing architectures and strategies for leveraging image-to-text transformations in sentiment prediction.
@@ III Datasets and Problem Formulation
In the present study, we use two datasets: PerceptSent [5] and DeepSent [26]. The former is a publicly available [^2] https://github.com/ceslop84/perceptsent. collection of 5,000 images, most (92.7%) of which depict outdoor scenes. The images have been sourced from Instagram, Flickr, and NYC311333https://portal.311.nyc.gov.. The dataset has been fully annotated by human evaluators using mainly Amazon Mechanical Turk [^4] https://www.mturk.com. (AMT). The latter consists of 1,269 images collected from X (old Twitter), from indoor and outdoor scenes, which have also been labeled through the AMT crowd-sourcing platform.
To align with prior work and facilitate reproducibility, we adopt the problem formulation for sentiment analysis proposed by Lopes [5]. Denoted here as $\langle{\sigma_{l}},{P}_{{C}}\rangle$, we have $C$ as the number of sentiment categories considered in a problem instance $P_{C}$, and $l$ as a threshold for filtering out divergent evaluators opinions, allowing us to define different types of sentiment dominance for filter $\sigma_{l}$.
@@
Fig. 1 presents one sample image from the PerceptSent dataset and one from the DeepSent dataset, as well as the data associated with each sample. Notably, in PerceptSent, the human-provided perceptions, along with the number of perceptions reported by each annotator, offer insight into the rationale behind each sentiment label. In contrast, for the DeepSent sample, the absence of such perceptions means we cannot infer why three annotators labeled the image as positive while two labeled it as negative.
Figure 1: Image samples and associated data from (a) PerceptSent and (b) DeepSent datasets, with the corresponding sentiment polarity vector ($\mathbf{s}$) assigned by human evaluators. Image source: [5, 26].
@@ IV-A Multimodal Large Language Models (MLLMs)
To ensure effective task alignment, we have experimented with the aforementioned MLLMs using various prompt formulations, including those generated by the models themselves. We selected the most effective prompt for each task, balancing clarity, consistency, and model response accuracy:
@@ IV-B Text-only Large Language Models (LLMs)
When adjusting the weights in Task 2a and Task 2b, the training loop iteratively updates model parameters over several epochs and incorporates early stopping based on validation F1-score improvements to prevent overfitting. The instances used as LLM inputs are textual descriptions provided by MLLMs. Such descriptions are framed as text generation tasks using sentiment-aware prompts adapted to the number of classes in each configuration $P_{C}$, where $C\in\{2,3,5\}$. Each prompt instructs the model to determine the sentiment of a given description by selecting from a predefined set of labels. For example, in the three-class setting $P_{3}$, the prompt is ”What is the sentiment of this description? Please choose an answer from {“Positive”: 2, “Negative”: 0, “Neutral”: 1}”. Then they are tokenized and go through data-loaders with dynamic batching and shuffling for robust training.
@@ V-A Direct Sentiment Classification using MLLMs
As shown in the third column of Table I (Task 1), the lowest $F$-scores for GPT (OAI) and DeepSeek occur under more challenging conditions, characterized by low annotator agreement, while higher scores are observed as annotator consensus increases. The same trend is observed with respect to the number of target classes; lower scores are obtained under harsher conditions, i.e., as the number of classes increases. GPT (OAI) achieves higher F-scores than DeepSeek, with relative gains [^5] For example, the relative $F$-score gain of GPT (OAI) regarding DeepSeek (DS) is calculated as (OAI - DS)/DS. of 2% to 67%. The improvement is statistically significant (paired t-test, $p<0.05$) in three cases; however, in the $\langle\sigma_{5},P_{3}\rangle$ configuration, the one with the lowest difference, this value is not significant ($p=0.194$).
In contrast, GPT (OS) fails to produce valid classification outputs for Task 1. As illustrated in Fig. 3, even after extensive prompt engineering — including explicit instructions such as ”Select a single sentiment class for this image from the $\{\mathcal{L}\}$ list”, where $\{\mathcal{L}\}$ refers to the list of sentiment polarities for a specific problem setup, or ”Do not describe this image; just choose one sentiment polarity from the $\{\mathcal{L}\}$ list” — the model consistently generates explanatory or ambiguous responses that attempt to justify the set $\{\mathcal{L}\}$ of categories, rather than selecting a single label, or combining multiple sentiment cues (e.g., ”Positive, but with some neutral elements”), making it difficult to unambiguously assign a single label and complicating automatic parsing. This behavior contrasts with the other MLLMs, which produce discrete and task-aligned outputs, as also shown in Fig. 3.
@@
Refer to caption Problem setup: P3 Dominance threshold: $\sigma=3$ Labeled as: Negative Task 1 GPT (OS): Positive: The image shows a group of people walking on the sidewalk, which suggests that they are out and about, enjoying their day or going about their daily business. The bright sunlight and clear blue sky add to the positive sentiment of the image. Neutral: The image is a reflection in a round mirror, making it difficult to determine the sentiment of the scene beyond the mirror … Negative: It is not possible to determine the sentiment of the image due to the reflection in the round mirror. The reflection obscures the view … GPT (OAI): Negative DeepSeek: Neutral 10121Evaluator’s votes
@@ V-B Pre-trained LLMs: classifying MLLMs image descriptions
V-C Fine-tuning LLMs for classifying MLLM image descriptions
Fine-tuning the LLMs (Task 2b in Table I) yields the highest overall $F$-scores. Relative gains over Task 2a range from [11.59%, 50.87%] for BART, [5.74%, 32.20%] for MBERT, and [4.09%, 102.87%] for LLAMA, across all MLLMs, evaluators’ agreement thresholds, and number of sentiment classes.
@@ V-C Fine-tuning LLMs for classifying MLLM image descriptions
Fig. 5 showcases four examples from the PerceptSent dataset alongside sentiment predictions from all considered models in both Task 2a and Task 2b.
@@
Figs. 5(a,b) present image samples with strong consensus ($\sigma_{5}$) among human evaluators. In Fig. 5(a), the image has received unanimous positive sentiment annotations from all five annotators. The associated human perceptions are strongly positive, with descriptions such as nature (mentioned four times) and pleasant environment (three times), among others. After fine-tuning, all MLLM+LLM combinations correctly predict the positive sentiment based on generated descriptions referencing visual elements such as a body of water, a beach, and sparkling highlights. Fig. 5(b) similarly shows high agreement, with all MLLM+LLM combinations (pre-trained and fine-tuned) accurately identifying the target negative sentiment. Human evaluators associated this sentiment with the presence of graffiti (noted by all annotators), lack of maintenance (three mentions), destruction (one mention), among other negative perceptions. The MLLMs image description results often capture these perceptions, highlighting features such as an underground tunnel, a graffiti-covered entrance, and a gritty atmosphere, with GPT (OAI) explicitly mentioning a sense of decay. Interestingly, both GPT (OAI) and DeepSeek can even recognize and transcribe the “OHNO” graffiti above the entrance. As shown in Table I, these two cases exemplify the best observed performance, where high evaluator agreement ($\sigma_{5}$) correlates with elevated $F$-scores. In contrast, under lower agreement ($\sigma_{3}$), some issues occur. Although in Fig. 5(c) most of the pre-trained models correctly classify slightly negative sentiment, some BART predictions shift toward wrong sentiment after fine-tuning. This change may result from the influence of image descriptions generated by GPT (OAI), such as “The atmosphere looks calm and residential, typical of a suburban area,” which potentially moderates the models’ sentiment evaluations. Nevertheless, fine-tuning improvements occur, particularly for GPT (OS)+LLAMA and DeepSeek+MBERT. In the case illustrated by Fig.5(d), none of the MLLM+LLM combinations succeed in predicting the correct negative sentiment label after fine-tuning (Task 2b). This fact appears to be driven by limitations in the scene descriptions generated by the MLLMs. Specifically, none of the models could identify salient visual features indicative of a negative context — such as the pile of trash on the left or the overall untidy condition of the street, as indicated in the human evaluator perceptions (marked as trash (two times), pollution and lack of maintenance). Instead, the generated image descriptions emphasize neutral or positive aspects, including references to a city street with parked cars, a calm atmosphere, and the presence of a few pedestrians. Such framing has likely influenced the LLMs to interpret the scene as neutral. Notably, LLAMA in Task 2a predicts the correct negative sentiment; however, with low $F$-scores ($30.3\%,46.6\%]$) for $\langle\sigma_{3},P_{3}\rangle$, this does not indicate consistent performance. After fine-tuning, the prediction shifted to neutral, confirming that the generated descriptions failed to capture the negative elements identified by human evaluators.
@@
Refer to caption Problem setup: P5 Dominance threshold: $\sigma_{5}$ Labeled as: Positive Human perceptions (# of votes): colors (1), leisure/fun/rest (1), sky (1), nature (4), pleasant environment (3). GPT (OS) GPT (OAI) DeepSeek BART Task 2a: Positive Positive Slightly Pos. Task 2b: Positive Positive Positive MBERT Task 2a: Slightly Pos. Positive Positive Task 2b: Positive Positive Positive LLAMA Task 2a: Positive Positive Positive Task 2b: Positive Positive Positive 50000Evaluator’s votes(a)
@@
Refer to caption Problem setup: P3 Dominance threshold: $\sigma_{5}$ Labeled as: Negative Human perceptions (# of votes): lack of maintenance (3), trash (1) graffiti (5), negative text message (1) debris/destruction (1), violence (1) GPT (OS) GPT (OAI) DeepSeek BART Task 2a: Negative Negative Negative Task 2b: Negative Negative Negative MBERT Task 2a: Negative Negative Negative Task 2b: Negative Negative Negative LLAMA Task 2a: Negative Negative Negative Task 2b: Negative Negative Negative 00014Evaluator’s votes(b)
@@
Refer to caption Problem setup: P5 Dominance threshold: $\sigma_{3}$ Labeled as: Slightly Negative Human perceptions (# of votes): trash (4), pollution (1) debris/destruction (1), lack of maintenance (1) animals (1), everyday image (1) GPT (OS) GPT (OAI) DeepSeek BART Task 2a: Slightly Neg. Slightly Neg. Slightly Neg. Task 2b: Negative Neutral Slightly Neg. MBERT Task 2a: Slightly Neg. Slightly Neg. Negative Task 2b: Slightly Neg. Slightly Neg. Slightly Neg. LLAMA Task 2a: Positive Neutral Slightly Neg. Task 2b: Slightly Neg. Neutral Slightly Neg. 00131Evaluator’s votes(c)
@@
Refer to caption Problem setup: P3 Dominance threshold: $\sigma_{3}$ Labeled as: Negative Human perceptions (# of votes): everyday image (1), accident (1) lack of colors (1), trash (2) pollution (1), lack of maintenance (1) GPT (OS) GPT (OAI) DeepSeek BART Task 2a: Positive Neutral Positive Task 2b: Neutral Neutral Positive MBERT Task 2a: Neutral Neutral Neutral Task 2b: Neutral Neutral Neutral LLAMA Task 2a: Negative Negative Negative Task 2b: Neutral Neutral Neutral 00212Evaluator’s votes(d)
@@ VI Conclusions and future work
Although we evaluated GPT-4o mini, MiniGPT-4 and DeepSeek-VL2-Tiny, our methodology is model-agnostic. The consistent performance patterns suggest newer MLLMs would yield further improvements while maintaining interpretability benefits. Notably, ModernBERT’s superior performance over larger models demonstrates that careful architecture selection and tuning can outweigh pure scale advantages. Moreover, despite GPT-4o mini achieving the best overall results, we note that open-weights (DeepSeek) or open source (GPT (OS)) alternatives can still be attractive compared to paid commercial models such as those from OpenAI. In that regard, we highlight that, due to GPU and time constraints, we employed the Tiny DeepSeek variant, and that larger models could achieve performance levels closer to that of GPT-based models. Thus, we believe a deeper exploration of open weights and open source models remains an interesting road for future research.
@@
Future research may investigate additional (and perhaps more powerful) MLLMs, and some may indeed surpass the results reported here. Yet such improvements would only strengthen our key finding: that robust multimodal language models are inherently well-suited for understanding sentiment polarity. Complementary avenues include (1) applying xAI approaches to illuminate the image regions and descriptive words that most influence sentiment predictions, and (2) developing interactive annotation interfaces that enable users to correct ambiguous image descriptions, ultimately improving dataset quality and model reliability.